3. Imagine a perfectly competitive firm producing good A with cost function TC=400+20Q-2Q 2 +2/3Q 3 , where Q is quantity produced

a. determine the firm’s short run supply curve

b. What is the profit maximizing level of output when price of A is birr 180?

4. Suppose the perfectly competitive price is given as $46 and the total cost of the firm is given by TC=14X+2X 2 , find

a. The profit maximizing level of output for the firm?

b. The profit of the firm?


1
Expert's answer
2021-12-21T12:08:14-0500

3.

a). Firm’s short run supply curve

A perfectly competitive firm producing good A with cost function;

"TC=400+20Q-2Q^2+\frac{2}{3}Q^3"

Firm's short run supply curve is marginal cost (MC).

"MC=\frac{dTC}{dQ}=\frac{d}{dQ}(400+20Q-2Q^2+\frac{2}{3}Q^3)"

"MC=20-4Q+2Q^2"

b). Profit maximizing level of output

In case of perfectly competitive firm, profit maximizing occurs when;

"MC=P"

"20-4Q+2Q^2=180"

"2Q^2-4Q-160=0"

"Q^2-2Q-80=0"

"Q^2-10Q+8Q-80=0"

"Q(Q-10)+8(Q-10)=0"

"Q-10=0"

"Q=10"

"Profit=(180\times10)-[20-(4\times10)+2(10)^2]"

"Profit=\$1620"


4.

a). Profit maximizing level of output

Price = $46

"TC=14X+2X^2"

"MC=\frac{dTC}{dQ}=\frac{d}{dQ}(14X+2X^2)"

"MC=X+4X"

"MC=P"

"X+4X=46"

"5X=46"

"X=\frac{46}{5}=9.2\approx9"

b). Profit of the firm

"Profit=(46\times9)-[9+(4\times9)]"

"Profit=\$369"